# Shane Cooper and Ilia Kamotski

### v1

@@ Abstract
We consider ‘off-axis’ electromagnetic wave propagation down the homogeneous direction of a low-loss two-dimensional periodic dielectric (or photonic crystal fibre) near the light line of the low-dielectric material constituent. Numerous physical and numerical experiments demonstrate the presence of photonic band gaps in the vicinity of this ‘critical’ light line. We mathematically analyse the existence of photonic band gaps near the line and characterise them in terms of frequency gaps in the spectrum of the Maxwell equations restricted to the line. We apply the results to both one-dimensional photonic crystal fibres, and a genuinely two-dimensional photonic crystal fibres with ‘thin’ inclusions, namely “ARROW” fibres. In the case of ARROW fibres, by an asymptotic analysis, in terms of the small inclusion parameter, we demonstrate the existence of low frequency photonic band gaps.
@@ 1 Introduction
A two-dimensional photonic crystal fibre is a low-loss periodic dielectric with material properties that are homogeneous along one of its axes. Photonic crystal fibres (PCF) are known to possess photonic band gaps; that is some PCF prevent certain frequency ranges of light from propagating in some, or all, directions. Such devices are of practical importance as the can be used to construct wave guides with optical performances that surpass traditional optical wave-guides (that work on the principle of total internal reflection), such as omni-directional reflectors and narrow-band filters, to name a few (there is too much literature to reference but refer to the book [6] and citations therein). Consequently, an important mathematical task is to analytically characterise photonic band gaps in terms of a PCFs material and geometric properties.
@@
In physical literature, the origin of a photonic band gap is explained as follows (cf. [10]): suppose we are considering plane-wave solutions to the Maxwell system with frequency $\omega$ and wave number $k$ down a given axis, through a given homogeneous dielectric material with constant dielectric permittivity $\epsilon$. Then, the $(\omega,k)$-plane is divided into two regions by the ‘light line’ $\omega(k)=ck/\sqrt{\epsilon}$, for $c$ the speed of light; above this line, waves propagate in the direction perpendicular to the axis, and below this line, waves are evanescent (i.e. decay) in the perpendicular direction. Now, for propagation down a photonic crystal fibre, with wave number $k$ and frequency $\omega$, the key observation of the above argument is this: if $(\omega,k)$ is above the light line for the low-dielectric matrix phase, the whole PCF can support the wave but just below the low-dielectric matrix light line (and above the high-dielectric inclusion light line) the matrix phase can no longer support propagating EM waves. It is argued that, for this reason, one should expect photonic band gaps to appear near the low-dielectric inclusion light line. This heuristic argument is supported in literature through numerous physical and numerical experiments ([10], [8] [11] to name a few).
@@
In this work we provide a rigorous mathematical analysis of the above heuristics. We henceforth refer to the low-dielectric inclusion light line as the critical (light) line. Before, discussing this article’s results, we review previous related mathematical results. For non-trivial wave propagation down the PCF, i.e. $k\neq 0$, then the EM field does not admit polarisations; consequently, it can been seen that EM propagation directly above the critical line reduces the Maxwell system to a system of two second-order elliptic differential spectral equations with coefficients that become highly contrasting and an ellipticity constant that degenerates to zero as one approaches the critical light line. In [4], the author developed further two-scale resolvent convergence methods, of [12], [13], and [7], to analyse the asymptotics of the Bloch spectrum of the associated degenerating high-contrast differential operator, as one approached the critical light line. It was shown the Bloch spectrum converged to the spectrum of some ‘limit’ operator and that photonic band gaps near the critical line will appear if the determined limit operator’s spectrum possessed gaps.
@@
In this work, we improve these results in several key ways. First, we establish that the limit operator, found in [4], is precisely the Maxwell operator restricted to the critical line. Furthermore, we utilise some methods of [5] to improve the above convergence results to provide operator-type error estimates between the Maxwell operators on and above the critical light line, respectively, in terms of the distance to the line. This implies, amongst other things, that if the Maxwell operator, on the critical light line, has gaps in its spectrum then photonic band gaps exist in a quantifiable neighbourhood of the restricted Maxwell spectral gaps. Finally, we study the particular problem of ARROW fibres, i.e. PCF with thin inclusions, to demonstrate the existence of low-frequency gaps in this restricted Maxwell spectrum and therefore, by the above results, prove the existence of low-frequency photonic band gaps for ARROW PCF.
@@
$\R.\phi_{1,1}+\phi_{2,2}=0\quad\text{and}\quad\phi_{2,1}-\phi_{1,2}=0\quad\text{in $\square\backslash R$.}$ (1.2)
@@
[…]
@@
$\sigma_{0}=\overline{\sum_{n=1}^{\infty}[a_{n},b_{n}]},\quad a_{n}:=\min_{\theta\in\square^{*}}\lambda_{n}(\theta),b_{n}:=\max_{\theta\in\square^{*}}\lambda_{n}(\theta),$
@@
$\lambda_{n}(\theta)=\frac{1}{\delta^{2}\big{(}g_{\theta}(0)-(2\pi)^{-1}\ln{\delta}\big{)}}\Lambda_{n}+\mathcal{O}(\delta^{-2}|\ln{\delta}|^{-3/2}),\quad n=1,2.$
@@
$\lambda_{n}(\theta)=\delta^{-2}\Lambda_{n}+\mathcal{O}(\delta^{-2}|\ln{\delta}|^{-1/2}),\quad n\geq 3,$
@@ 2 Problem formulation
[…]
@@
×$\nabla\times\tilde{H}$ $=-\,\epsilon\,\tfrac{\partial\tilde{E}}{\partial t},$ (2.3)
×$\nabla\times\tilde{E}$ $=\,\tfrac{\partial\tilde{H}}{\partial t}.$
@@
$E(y+e_{j})=e^{\mathrm{i}\theta_{j}}E(y),\quad\&\quad H(y+e_{j})=e^{\mathrm{i}\theta_{j}}H(y),\quad y\in\square,\,j=1,2,$ (2.4)
@@
$\mathrm{div}\,u:=u_{1,1}+u_{2,2}\quad\text{and}\quad\mathrm{curl}\,u:=u_{2,1}-u_{1,2}$
@@
$\mathrm{div}\,u=0\quad\text{and}\quad\mathrm{curl}\,u=0\quad\text{in $\square\backslash R$,}$ (2.7)
@@
$b(u,\phi)=\lambda\,c(u,\phi),\qquad\text{where\ }\lambda=\omega^{2},$ (2.8)
@@
$b(u,\phi)$ $:=-\frac{1}{2}\int_{\square\backslash R}\Big{(}\nabla^{\perp}u_{1}\cdot\overline{\nabla\phi_{2}}+\nabla u_{2}\cdot\overline{\nabla^{\perp}\phi_{1}}\Big{)}+$ (2.9)
@@
•  In \$\𝑅\square\backslash R,$ we notice that the matrices in (2.5) have non-trivial kernels. This provides constraints on $E_{3},H_{3}$. Indeed, we see that
$\R\left(\begin{array}[]{cc}k&-\epsilon\\ -1&k\end{array}\right)=\left(\begin{array}[]{cc}1&-1\\ -1&1\end{array}\right)\quad\text{in $\square\backslash R$}$
@@
$\R\left(\begin{array}[]{c}H_{1}\\ E_{2}\end{array}\right)=-\frac{\mathrm{i}}{\omega}\frac{1}{2}\left(\begin{array}[]{c}H_{3,1}\\ E_{3,2}\end{array}\right)+C_{1}\left(\begin{array}[]{c}1\\ 1\end{array}\right)\quad\text{in $\square\backslash R$}$
@@
$\R\left(\begin{array}[]{c}H_{2}\\ E_{1}\end{array}\right)=-\frac{\mathrm{i}}{\omega}\frac{1}{2}\left(\begin{array}[]{c}H_{3,2}\\ E_{3,1}\end{array}\right)+C_{2}\left(\begin{array}[]{c}-1\\ 1\end{array}\right)\quad\text{in $\square\backslash R$}$
@@
$\mathrm{div}\,\phi=0,\quad\mathrm{curl}\,\phi=0,\quad\text{in $\square\backslash R$,}$
@@
\$\int_{\square\backslash R}\left(\begin{array}[]{c}H_{1}\\ E_{2}\end{array}\right)\cdot\overline{\left(\begin{array}[]{c}\phi_{2,2}\\ \phi_{1,1}\end{array}\right)}-\left(\begin{array}[]{c}H_{2}\\ E_{1}\end{array}\right)\cdot\overline{\left(\begin{array}[]{c}\phi_{2,1}\\ \phi_{1,2}\end{array}\right)}=-\frac{\mathrm{i}}{\omega}\frac{1}{2}\int_{\square\backslash R}\Big{(}\nabla^{\perp}u_{1}\cdot\overline{\nabla\phi_{2}}+\nabla u_{2}\cdot\overline{\nabla^{\perp}\phi_{1}}\Big{)}$
@@
$\left(\begin{array}[]{c}H_{1}\\ E_{2}\end{array}\right)=\frac{\mathrm{i}}{\omega}\frac{1}{\gamma}\left(\begin{array}[]{c}H_{3,1}+\epsilon_{0}E_{3,2}\\ H_{3,1}+E_{3,2}\end{array}\right),\quad\left(\begin{array}[]{c}H_{2}\\ E_{1}\end{array}\right)=\frac{\mathrm{i}}{\omega}\frac{1}{\gamma}\left(\begin{array}[]{c}H_{3,2}-\epsilon_{0}E_{3,1}\\ E_{3,1}-H_{3,2}\end{array}\right),\quad\text{in $R$}.$
@@
Here, we consider $k=\sqrt{1-\varepsilon}$ for fixed $10<\varepsilon<$ 1. We shall show that the problem (2.4)-(2.6) is equivalent to $\theta$-quasi-periodic $u=(H_{3},E_{3})$ and $\lambda=\omega^{2}$ satisfying
@@
$a(u,\phi):=\int_{\square\backslash R}\mathrm{div}\,u\,\overline{\mathrm{div}\,\phi}+\mathrm{curl}\,{u}\,\overline{\mathrm{curl}\,\phi},$ (2.12)
$b_{\varepsilon}(u,\phi):=\int_{\square\backslash R}\frac{\sqrt{1-\varepsilon}-1}{\varepsilon}\Big{(}\nabla^{\perp}u_{1}\cdot\overline{\nabla\phi_{2}}+\nabla u_{2}\cdot\overline{\nabla^{\perp}\phi_{1}}\Big{)}\ +$ (2.13)
@@
$A\,\,:=\left\{\begin{matrix}\varepsilon&&\text{in}\ \square\backslash R,\\ \gamma+\varepsilon&&\text{in}\ R,\end{matrix}\right.$
@@
$V(\theta):=\{v\in H^{1}_{\theta}(\square;\mathbb{C}^{2})\ |\ \mathrm{div}\,v=0\quad\&\quad\mathrm{curl}\,v=0\ \text{in $\square\backslash R$}\},\quad\theta\in\square^{*}.$ (2.17)
@@
$b(v,\tilde{v}):=\int_{\square}\nabla v_{2}\cdot\overline{\nabla\tilde{v}_{2}}+\gamma^{-1}\int_{\square}\big{(}\mathrm{div}\,{v}\,\overline{\mathrm{div}\,{\tilde{v}}}+\mathrm{curl}\,{v}\,\overline{\mathrm{curl}\,{\tilde{v}}}\big{)},\quad\forall v,\tilde{v}\in V(\theta).$ (2.18)
@@
$\int_{\square}(\mathrm{div}\,u\,\overline{\mathrm{div}\,\phi}+\mathrm{curl}\,{u}\,\overline{\mathrm{curl}\,\phi})=\int_{\square}\nabla u:\overline{\nabla\phi},\qquad\forall u,\phi\in H^{1}_{\theta}(\square;\mathbb{C}^{2}).$ (2.19)
@@
$b(v,\tilde{v})$ $=\int_{\square}\nabla v_{2}\cdot\overline{\nabla\tilde{v}_{2}}+\gamma^{-1}\nabla v:\overline{\nabla\tilde{v}},\quad\forall v,\tilde{v}\in V(\theta).$ (2.20)
@@
:=$\R\mathcal{H}:=\{h\in L^{2}(\square;\mathbb{C}^{2})\,|\,\mathrm{div}\,{h}=\mathrm{curl}\,{h}=0\text{ in $\square\backslash R$}\},$
@@
$b(v,\tilde{v})=c(\mathcal{B}(\theta)v,\tilde{v}),\quad\forall\,v\in\mathrm{dom}(\mathcal{B}(\theta)),\,\forall\,\tilde{v}\in V(\theta).$
@@
$\varepsilon^{-1}a(\phi,\psi)+b_{\varepsilon}(\phi,\psi)=c(\mathcal{B}_{\varepsilon}(\theta)\phi,\psi),\quad\forall\,\phi\in\mathrm{dom}(\mathcal{B}_{\varepsilon}(\theta)),\,\forall\,\psi\in H^{1}_{\theta}(\square;\mathbb{C}^{2}).$
@@
As $\mathcal{B}_{\varepsilon}(\theta)$ has a compact resolvent, then non-trivial waves propagate down the two-dimensional photonic crystal fibre, for $k=\sqrt{1-\varepsilon}$, $10<\varepsilon<$ 1, if, and only if, $\lambda=\omega^{2}$ is an eigenvalue of $\mathcal{B}_{\varepsilon}(\theta)$ for some $\theta\in\square^{*}$.
@@
\$\int_{\square\backslash R}\left(|\mathrm{div}\,\phi|^{2}+|\mathrm{curl}\,\phi|^{2}+(\sqrt{1-\varepsilon}-1)\big{(}\nabla^{\perp}\phi_{1}\cdot\overline{\nabla\phi_{2}}+\nabla\phi_{2}\cdot\overline{\nabla^{\perp}\phi_{1}}\big{)}\ \right)$
=\$=\int_{\square\backslash R}\left(|\nabla\phi|^{2}+\sqrt{1-\varepsilon}\big{(}\nabla^{\perp}\phi_{1}\cdot\overline{\nabla\phi_{2}}+\nabla\phi_{2}\cdot\overline{\nabla^{\perp}\phi_{1}}\big{)}\ \right)$
\$\geq\int_{\square\backslash R}\left(|\nabla\phi|^{2}-2\sqrt{1-\varepsilon}|\nabla\phi_{1}||\nabla\phi_{2}|\right)$
\$\geq\int_{\square\backslash R}(1-\sqrt{1-\varepsilon})|\nabla\phi|^{2}$
@@
Similarly, for $10<\eta<$ 1 the positive solution to $1-\eta=\epsilon_{0}-\eta^{-1}$, one has
@@
$\varepsilon^{-1}a[\phi]+b_{\varepsilon}[\phi]\geq\frac{1}{2}\|\nabla\phi\|^{2}_{L^{2}(\square\backslash R;\mathbb{C}^{2})}+\frac{1}{\gamma+\varepsilon}(1-\eta)\|\nabla\phi\|^{2}_{L^{2}(R;\mathbb{C}^{2})}\geq\nu\|\nabla\phi\|^{2}_{L^{2}(\square;\mathbb{C}^{2})}$
@@
$\varepsilon^{-1}a(U_{\varepsilon,\theta},\phi)+b_{\varepsilon}(U_{\varepsilon,\theta},\phi)+c(U_{\varepsilon,\theta},\phi)=\langle F,\phi\rangle,\quad\forall\phi\in H^{1}_{\theta}(\square;\mathbb{C}^{2}),$
@@
$b(V_{\theta},\tilde{V})+c(V_{\theta},\tilde{V})=\langle F,\tilde{V}\rangle,\quad\forall\,\tilde{V}\in V(\theta).$
@@
$\|(\mathcal{B}_{\varepsilon}(\theta)+I)^{-1}-(\mathcal{B}(\theta)+I)^{-1}P\|_{L^{2}(\square;\mathbb{C}^{2})\rightarrow L^{2}(\square;\mathbb{C}^{2})}\leq C\varepsilon,\quad\forall\theta\in\square^{*},\forall 0<\varepsilon<1.$
Lemma 3.2, and the fact that $V(\theta)$ is infinite-dimensional, readily provides a uniform asymptotic comparision on spectra(see [5, Section 6] for details). Indeed, let
$0\leq\lambda_{1,\varepsilon}(\theta)\leq\lambda_{2,\varepsilon}(\theta)\leq\lambda_{3,\varepsilon}(\theta)\leq\ldots,\qquad\lambda_{n,\varepsilon}(\theta)\rightarrow\infty\ \text{as}\ n\rightarrow\infty.$
@@
$0\leq\lambda_{1}(\theta)\leq\lambda_{2}(\theta)\leq\lambda_{3}(\theta)\leq\ldots,\qquad\lambda_{n}(\theta)\rightarrow\infty\ \text{as}\ n\rightarrow\infty.$
@@
[…]
@@
Indeed, the following result holds (see, for example, [5, Section 6] for details):
@@
$\sigma_{0}=\overline{\sum_{n=1}^{\infty}[a_{n},b_{n}]}\quad a_{n}:=\min_{\theta\in\square^{*}}\lambda_{n}(\theta),b_{n}:=\max_{\theta\in\square^{*}}\lambda_{n}(\theta),$

### v2

@@ Abstract
We consider ‘off-axis’ electromagnetic wave propagation down the homogeneous direction of a low-loss two-dimensional periodic dielectric (or photonic crystal fibre) near the light line of the low-dielectric material constituent. Numerous physical and numerical experiments demonstrate the presence of photonic band gaps in the vicinity of this ‘critical’ light line. We mathematically analyse the existence of photonic band gaps near the line and characterise them in terms of frequency gaps in the spectrum of the Maxwell equations restricted to the line. We apply the results to both one-dimensional photonic crystal fibres, and a genuinely two-dimensional photonic crystal fibres with ‘thin’ inclusions, namely “ARROW” fibres. In the case of ARROW fibres, by an asymptotic analysis, in terms of the small inclusion parameter, we demonstrate the existence of low frequency photonic band gaps. It is important to note that our analysis does not assume any specific ratio between the dielectric contrasts; as a result, moderate or even low contrast models are fully encompassed. Similarly, no limiting assumptions are imposed on the wave propagation constant along the fibre.
@@ 1 Introduction
A two-dimensional photonic crystal fibre is a low-loss periodic dielectric with material properties that are homogeneous along one of its axes. Photonic crystal fibres (PCFs) are known to possess photonic band gaps; that is, some PCFs prevent certain frequency ranges of light from propagating in some, or all, directions. Such devices are of practical importance, as they can be used to construct waveguides with optical performances that surpass traditional optical waveguides (which operate on the principle of total internal reflection), such as omnidirectional reflectors and narrow-band filters (see [10] and references therein). Consequently, an important mathematical task is to analytically characterise photonic band gaps in terms of a PCF’s material and geometric properties.
@@
In physical literature, the origin of a photonic band gap is explained as follows (cf. [15]): suppose we consider plane-wave solutions to the Maxwell system with frequency $\omega$ and wave number $k$ down a given axis, through a given homogeneous dielectric material with constant dielectric permittivity $\epsilon$. Then, the $(\omega,k)$-plane is divided into two regions by the ‘light line’ $\omega(k)=ck/\sqrt{\epsilon}$, for $c$ the speed of light; above this line, waves propagate in the direction perpendicular to the axis, and below this line, waves are evanescent in the perpendicular direction. For propagation down a photonic crystal fibre, with wave number $k$ and frequency $\omega$, the key observation of the above argument is this: if $(\omega,k)$ is above the light line for the low-dielectric matrix phase, the whole PCF can support the wave, but just below the low-dielectric matrix light line (and above the high-dielectric inclusion light line) the matrix phase can no longer support propagating EM waves. It is argued that, for this reason, one should expect photonic band gaps to appear near the low-dielectric inclusion light line. This heuristic argument is supported in literature through numerous physical and numerical experiments ([15], [12], [16]). We henceforth refer to the low-dielectric inclusion light line as the critical (light) line.
In this work we provide a rigorous quantitative mathematical analysis of the above heuristics. Before discussing this article’s results, we review previous related mathematical results. The rigorous mathematical theory of band gaps has largely focused on trivial wave number $k$ regimes or the high-contrast between material parameters regime. In classical theory, $k=0$ is assumed and the Maxwell system decouples to scalar Helmholtz equation due to high symmetries that allow polarisation. In the setting $k\rightarrow\infty$ the Maxwell system asymptotically decouples once more to the scalar regime and this was rigorously analysed in [2] via asymptotic methods. As for the high-contrast regime, various asymptotic approaches have been used to yield explicit spectral gap results (e.g., [9]; [1], [3], [18]) and the [14] considered how large the contrast must be to induce spectral gaps.
@@
In this work, we do not assume the high-contrast regime; in fact, for non-trivial wave propagation down the photonic crystal fibre, i.e., $k\neq 0$ nor $k\not\rightarrow\infty$ , the EM field does not admit polarisations; consequently, EM propagation directly above the critical line reduces the Maxwell system to a system of two second-order elliptic differential spectral equations with coefficients that artificially” become highly contrasting and an ellipticity constant that degenerates to zero as one approaches the critical light line ($k\rightarrow 1$). In [7], the author developed further two-scale resolvent convergence methods, of [17], [18], and [11], to analyse the asymptotics of the Bloch spectrum of the associated degenerating high-contrast differential operator, as one approached the critical light line. It was shown that the Bloch spectrum converged to the spectrum of some ‘limit’ operator, and that photonic band gaps near the critical line will appear if the determined limit operator’s spectrum possessed gaps.
@@
Here, we improve on the results of [11] in several key ways. First, we establish that the limit operator, found in [7], is precisely the Maxwell operator restricted to the critical line. Furthermore, we utilise some methods of [8] to improve the above convergence results to provide operator-type error estimates between the Maxwell operators on and above the critical light line, respectively, in terms of the distance to the line. This implies, among other things, that if the Maxwell operator on the critical light line has gaps in its spectrum, then photonic band gaps exist in a quantifiable neighbourhood of the restricted Maxwell spectral gaps. Finally, we study the particular problem of ARROW fibres, i.e., PCFs with thin inclusions, to demonstrate the existence of low-frequency gaps in this restricted Maxwell spectrum and therefore, by the above results, prove the existence of low-frequency photonic band gaps for ARROW PCF.
@@
[…]
@@
$\|(\mathcal{B}_{\varepsilon}(\theta)+I)^{-1}-(\mathcal{B}(\theta)+I)^{-1}P\|_{L^{2}(\square;\mathbb{C}^{2})\rightarrow L^{2}(\square;\mathbb{C}^{2})}\leq C\varepsilon,\qquad\forall\theta\in\square^{*},\forall\,0<\varepsilon<1.$
@@
$\sigma_{0}=\overline{\sum_{n=1}^{\infty}[a_{n},b_{n}]},\qquad a_{n}:=\min_{\theta\in\square^{*}}\lambda_{n}(\theta),b_{n}:=\max_{\theta\in\square^{*}}\lambda_{n}(\theta),$
@@
$\lambda_{n}(\theta)=\frac{1}{\delta^{2}\big(g_{\theta}(0)-(2\pi)^{-1}\ln{\delta}\big)}\Lambda_{n}+\mathcal{O}(\delta^{-2}|\ln{\delta}|^{-3/2}),\qquad n=1,2.$
@@
$\lambda_{n}(\theta)=\delta^{-2}\Lambda_{n}+\mathcal{O}(\delta^{-2}|\ln{\delta}|^{-1/2}),\qquad n\geq 3,$
@@ 2 Problem formulation
$\epsilon\big(x_{1},x_{2},x_{3}\big)=\left\{\begin{matrix}\epsilon_{0}&&\text{if $(x_{1},x_{2})\in\bigcup_{z\in\mathbb{Z}^{2}}(R+z)$,}\\[4.25006pt] 1&&\text{otherwise,}\end{matrix}\right$. (2.1)
@@
$\nabla\times\tilde{H}$ $=-\,\epsilon\,\tfrac{\partial\tilde{E}}{\partial t},$ (2.3)
$\nabla\times\tilde{E}$ $=\,\tfrac{\partial\tilde{H}}{\partial t}.$
@@
$E(y+e_{j})=e^{\mathrm{i}\theta_{j}}E(y),\qquad\&\qquad H(y+e_{j})=e^{\mathrm{i}\theta_{j}}H(y),\qquad y\in\square,\,j=1,2,$ (2.4)
@@
$\mathrm{div}\,u:=u_{1,1}+u_{2,2}\qquad\text{and}\qquad\mathrm{curl}\,u:=u_{2,1}-u_{1,2}$
@@
[…]
@@
$b(u,\phi)=\lambda\,c(u,\phi),\hskip 17.00024pt\text{where\ }\lambda=\omega^{2},$ (2.8)
@@
$b(u,\phi)$ $=-\frac{1}{2}\int_{\square\backslash R}\Big(\nabla^{\perp}u_{1}\cdot\overline{\nabla\phi_{2}}+\nabla u_{2}\cdot\overline{\nabla^{\perp}\phi_{1}}\Big)+$ (2.9)
@@
[…]
$\left(\begin{array}[]{cc}k&-\epsilon\\ -1&k\end{array}\right)=\left(\begin{array}[]{cc}1&-1\\ -1&1\end{array}\right)\qquad\text{in $\square\backslash R$}$
@@
$\left(\begin{array}[]{c}H_{1}\\ E_{2}\end{array}\right)=-\frac{\mathrm{i}}{\omega}\frac{1}{2}\left(\begin{array}[]{c}H_{3,1}\\ E_{3,2}\end{array}\right)+C_{1}\left(\begin{array}[]{c}1\\ 1\end{array}\right)\qquad\text{in $\square\backslash R$}$
@@
$\left(\begin{array}[]{c}H_{2}\\ E_{1}\end{array}\right)=-\frac{\mathrm{i}}{\omega}\frac{1}{2}\left(\begin{array}[]{c}H_{3,2}\\ E_{3,1}\end{array}\right)+C_{2}\left(\begin{array}[]{c}-1\\ 1\end{array}\right)\qquad\text{in $\square\backslash R$}$
@@
[…]
@@
$\int_{\square\backslash R}\left(\begin{array}[]{c}H_{1}\\ E_{2}\end{array}\right)\cdot\overline{\left(\begin{array}[]{c}\phi_{2,2}\\ \phi_{1,1}\end{array}\right)}-\left(\begin{array}[]{c}H_{2}\\ E_{1}\end{array}\right)\cdot\overline{\left(\begin{array}[]{c}\phi_{2,1}\\ \phi_{1,2}\end{array}\right)}=-\frac{\mathrm{i}}{\omega}\frac{1}{2}\int_{\square\backslash R}\Big(\nabla^{\perp}u_{1}\cdot\overline{\nabla\phi_{2}}+\nabla u_{2}\cdot\overline{\nabla^{\perp}\phi_{1}}\Big)$
@@
$\left(\begin{array}[]{c}H_{1}\\ E_{2}\end{array}\right)=\frac{\mathrm{i}}{\omega}\frac{1}{\gamma}\left(\begin{array}[]{c}H_{3,1}+\epsilon_{0}E_{3,2}\\ H_{3,1}+E_{3,2}\end{array}\right),\qquad\left(\begin{array}[]{c}H_{2}\\ E_{1}\end{array}\right)=\frac{\mathrm{i}}{\omega}\frac{1}{\gamma}\left(\begin{array}[]{c}H_{3,2}-\epsilon_{0}E_{3,1}\\ E_{3,1}-H_{3,2}\end{array}\right),\qquad\text{in $R$}.$
@@
Here, we consider $k=\sqrt{1-\varepsilon}$ for fixed $0<\varepsilon<1$. We shall show that the problem (2.4)-(2.6) is equivalent to $\theta$-quasi-periodic $u=(H_{3},E_{3})$ and $\lambda=\omega^{2}$ satisfying
@@
$a(u,\phi)=\int_{\square\backslash R}\mathrm{div}\,u\,\overline{\mathrm{div}\,\phi}+\mathrm{curl}\,{u}\,\overline{\mathrm{curl}\,\phi},$ (2.12)
$b_{\varepsilon}(u,\phi)=\int_{\square\backslash R}\frac{\sqrt{1-\varepsilon}-1}{\varepsilon}\Big(\nabla^{\perp}u_{1}\cdot\overline{\nabla\phi_{2}}+\nabla u_{2}\cdot\overline{\nabla^{\perp}\phi_{1}}\Big)\ +$ (2.13)
@@
[…]
@@
$V(\theta):=\{v\in H^{1}_{\theta}(\square;\mathbb{C}^{2})\ |\ \mathrm{div}\,v=0\qquad\&\qquad\mathrm{curl}\,v=0\ \text{in $\square\backslash R$}\},\qquad\theta\in\square^{*}.$ (2.17)
@@
$b(v,\tilde{v}):=\int_{\square}\nabla v_{2}\cdot\overline{\nabla\tilde{v}_{2}}+\gamma^{-1}\int_{\square}\big(\mathrm{div}\,{v}\,\overline{\mathrm{div}\,{\tilde{v}}}+\mathrm{curl}\,{v}\,\overline{\mathrm{curl}\,{\tilde{v}}}\big),\qquad\forall v,\tilde{v}\in V(\theta).$ (2.18)
@@
$\int_{\square}(\mathrm{div}\,u\,\overline{\mathrm{div}\,\phi}+\mathrm{curl}\,{u}\,\overline{\mathrm{curl}\,\phi})=\int_{\square}\nabla u:\overline{\nabla\phi},\hskip 17.00024pt\forall u,\phi\in H^{1}_{\theta}(\square;\mathbb{C}^{2}).$ (2.19)
@@
[…]
@@
$\mathcal{H}:=\{h\in L^{2}(\square;\mathbb{C}^{2})\,|\,\mathrm{div}\,{h}=\mathrm{curl}\,{h}=0\text{ in $\square\backslash R$}\},$
@@
$b(v,\tilde{v})=c(\mathcal{B}(\theta)v,\tilde{v}),\qquad\forall\,v\in\mathrm{dom}(\mathcal{B}(\theta)),\,\forall\,\tilde{v}\in V(\theta).$
@@
$\varepsilon^{-1}a(\phi,\psi)+b_{\varepsilon}(\phi,\psi)=c(\mathcal{B}_{\varepsilon}(\theta)\phi,\psi),\qquad\forall\,\phi\in\mathrm{dom}(\mathcal{B}_{\varepsilon}(\theta)),\,\forall\,\psi\in H^{1}_{\theta}(\square;\mathbb{C}^{2}).$
@@
As $\mathcal{B}_{\varepsilon}(\theta)$ has a compact resolvent, then non-trivial waves propagate down the two-dimensional photonic crystal fibre, for $k=\sqrt{1-\varepsilon}$, $0<\varepsilon<1$, if, and only if, $\lambda=\omega^{2}$ is an eigenvalue of $\mathcal{B}_{\varepsilon}(\theta)$ for some $\theta\in\square^{*}$.
@@
$\int_{\square\backslash R}\left(|\mathrm{div}\,\phi|^{2}+|\mathrm{curl}\,\phi|^{2}+(\sqrt{1-\varepsilon}-1)\big(\nabla^{\perp}\phi_{1}\cdot\overline{\nabla\phi_{2}}+\nabla\phi_{2}\cdot\overline{\nabla^{\perp}\phi_{1}}\big)\ \right)$
$=\int_{\square\backslash R}\left(|\nabla\phi|^{2}+\sqrt{1-\varepsilon}\big(\nabla^{\perp}\phi_{1}\cdot\overline{\nabla\phi_{2}}+\nabla\phi_{2}\cdot\overline{\nabla^{\perp}\phi_{1}}\big)\ \right)$
$\geq\int_{\square\backslash R}\left(|\nabla\phi|^{2}-2\sqrt{1-\varepsilon}|\nabla\phi_{1}||\nabla\phi_{2}|\right)$
[…]
@@
Similarly, for $0<\eta<1$ the positive solution to $1-\eta=\epsilon_{0}-\eta^{-1}$, one has
@@
[…]
@@
$\varepsilon^{-1}a(U_{\varepsilon,\theta},\phi)+b_{\varepsilon}(U_{\varepsilon,\theta},\phi)+c(U_{\varepsilon,\theta},\phi)=\langle F,\phi\rangle,\qquad\forall\phi\in H^{1}_{\theta}(\square;\mathbb{C}^{2}),$
@@
$b(V_{\theta},\tilde{V})+c(V_{\theta},\tilde{V})=\langle F,\tilde{V}\rangle,\qquad\forall\,\tilde{V}\in V(\theta).$
@@
$\|(\mathcal{B}_{\varepsilon}(\theta)+I)^{-1}-(\mathcal{B}(\theta)+I)^{-1}P\|_{L^{2}(\square;\mathbb{C}^{2})\rightarrow L^{2}(\square;\mathbb{C}^{2})}\leq C\varepsilon,\qquad\forall\theta\in\square^{*},\forall 0<\varepsilon<1.$
Lemma 3.2, and the fact that $V(\theta)$ is infinite-dimensional, readily provides a uniform asymptotic comparision on spectra(see [8, Section 6] for details). Indeed, let
$0\leq\lambda_{1,\varepsilon}(\theta)\leq\lambda_{2,\varepsilon}(\theta)\leq\lambda_{3,\varepsilon}(\theta)\leq\ldots,\hskip 17.00024pt\lambda_{n,\varepsilon}(\theta)\rightarrow\infty\ \text{as}\ n\rightarrow\infty.$
@@
$0\leq\lambda_{1}(\theta)\leq\lambda_{2}(\theta)\leq\lambda_{3}(\theta)\leq\ldots,\hskip 17.00024pt\lambda_{n}(\theta)\rightarrow\infty\ \text{as}\ n\rightarrow\infty.$
@@
$\left|\frac{1}{\lambda_{n,\varepsilon}(\theta)+1}-\frac{1}{\lambda_{n}(\theta)+1}\right|\leq C\varepsilon,\hskip 17.00024pt\forall\,n\in\mathbb{N},\forall\,\theta\in\square^{*},\forall\,0<\varepsilon<1.$
@@
Indeed, the following result holds (see, for example, [8, Section 6] for details):
@@
$\sigma_{0}=\overline{\sum_{n=1}^{\infty}[a_{n},b_{n}]}\qquad a_{n}:=\min_{\theta\in\square^{*}}\lambda_{n}(\theta),b_{n}:=\max_{\theta\in\square^{*}}\lambda_{n}(\theta),$